OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <10, 0>, v = <0, -9>

OpenStudy (anonymous):

Do you know about the dot product?

OpenStudy (anonymous):

Yea Sorry i was in the restroom :o

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Then what do you get when you do the dot product?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Okay, so what does it mean if the dot product between two vectors is 0?

OpenStudy (anonymous):

Is it parallel?

OpenStudy (anonymous):

<0,1> and <0,1> are identical vectors... clearly they are parallel. The dot product is 1. Thus I don't think it means that.

OpenStudy (anonymous):

so neither since there is no dot product D:

OpenStudy (anonymous):

No, the remaining options are orthogonal and neither.

OpenStudy (anonymous):

so its orthogonal if it isnt neither?

OpenStudy (anonymous):

Im sorry i just dont know how to find these :(

OpenStudy (anonymous):

It is either orthogonal or it is neither, but we need to find out first.

OpenStudy (anonymous):

Can you think of two vectors which are orthogonal? Do you know what orthogonal means?

OpenStudy (anonymous):

perpendicular

OpenStudy (anonymous):

okay, so can you think of two perpendicular vectors?

OpenStudy (anonymous):

i dont know :/

OpenStudy (anonymous):

Perpendicular means the angle between to vectors is \(90^\circ\). The dot product is equal to \(\|\mathbf v_1\|\|\mathbf v_2\|\cos\theta\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Where \(\theta\) is the angle between the vectors. Is this familiar at all?

OpenStudy (anonymous):

Yea alittle

OpenStudy (anonymous):

So what is \(\cos 90^\circ\)?

OpenStudy (anonymous):

ohhhhh Cos90=0

OpenStudy (anonymous):

so when the dot product is 0 it is perpendicular

OpenStudy (anonymous):

This is something that is supposed to be obvious.

OpenStudy (anonymous):

Well i apologize then. Can you help me with one more?

OpenStudy (anonymous):

Let u = <-6, -2>, v = <-2, 3>. Find -3u + 2v.

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

okay do you know how to multiply a vector by a scalar?

OpenStudy (anonymous):

i believe what i was told to do it -2(-6+-2) 2(-2,3) is that right?

OpenStudy (anonymous):

No, that is not correct. The answer will be a vector, not a number.

OpenStudy (anonymous):

oh then no i don't know

OpenStudy (anonymous):

Can you explain how to do it? :o

OpenStudy (anonymous):

Okay, so first what is \(-3\mathbf u\) equal to?

OpenStudy (anonymous):

Hint: it is another vector.

OpenStudy (anonymous):

\(-3\) is a scalar and \(\mathbf u\) is a vector... what happens when you multiply them?